I need help with this, please!!
The equation l(r) = 400r expresses the amount of simple interest earned by an investment of $400 for 1 year as a function of the rate of interest (r). Compute l(0.11), l(0.12), and l(0.15).

Answer by Earlsdon(6294)

If l(r) = 400r, find: l(0.11), l(0.12), and l(0.15)
Replace the r in the function with the values given in the parentheses.
l(0.11) = 400(0.11) = 44
l(0.12) = 400(0.12) = 48
l(0.15) = 400(0.15) = 60